OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? A. (1, 7) B. (5, 0) C. (2, -1) D. (0, 5)

OpenStudy (apoorvk):

Substitute the value of 'y' (in terms of 'x') from the first equation in the second equation. so, -3x+5 = 2x+5 solving this, we get, x=0 so subtitute this value of 'x' in the first equation, you get: y=-3(0)+5 or, y=5 hence x=0, y=-5 is a solution for the above system. Helps?

OpenStudy (apoorvk):

Or you could try @aroub 's method. That is checking every solution. But since this question has pretty easy figures, solving it methodically would require fractionally less time. Also, in cases where you don't have options, you'll have to use the 'snail' method.

OpenStudy (anonymous):

Okay, and thank you

OpenStudy (apoorvk):

lo, there is no 'Aroub's method" anymore - what she did was basically substituted every set of value in the set of equations, and checked which satisfies them - helps if you suddenly forgot the procedures (which I hope will not) and are at sea.

OpenStudy (aroub):

I think apoorvk's method is better :D So just stick to it. I do not want you to be confused.

OpenStudy (aroub):

The snail method? lol. which is?

OpenStudy (apoorvk):

The normal methods are all 'snail' methods - which one has no option but to use 'at times'.